OpenStudy (anonymous):

how to solve y=1/2x²-3x+3 using completing the square . ?

OpenStudy (shinalcantara):

is the equation\[y = \frac{ 1 }{ 2x ^{2} } - 3x + 3\] or \[y = \frac{ 1 }{ 2x ^{2} -3x + 3}\]